OpenStudy (anonymous):

The table below represents a linear function f(x) and the equation represents a function g(x): -1 -9 0 -1 1 7 g(x) g(x) = 3x - 2 Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x).

OpenStudy (anonymous):

i dont understand how to change that table into an equation. i know that the last numbers of the equation is - 1 because 0 times any number is 0 and then you get -1 because of you subtract 1

OpenStudy (anonymous):

To convert any table of data that represents only "linear function" then you have to know the general formula of any line which is known to be : \[y/f(x) = mx + c\] where m is the slope of that line while c is known as y-intercept .. the part that the line intercepts y-axis when x = 0 So now it is obvious that f(x) has the form mx+c

OpenStudy (triciaal):

a linear equation is of the form y = mx + b where m is the slope and b is the y value when x = 0 in the form of a table you are given the x and corresponding y value

OpenStudy (anonymous):

so do i plug them into the equation and solve?

OpenStudy (triciaal):

the slope of a line is the change in y divided by the corresponding change in x

OpenStudy (triciaal):